# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v2

@@ B.2 Lower bound $\widetilde{\alpha}_{0}(\mathbf{x})$
${\sf acc}({\widehat{\bm{\theta}}}_{0})\geq 1-\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\gamma^{4}}{16\lambda_{\max}^{2}d}\right)-e^{-d/8}-2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$
@@
Proof Since we are interested with dependence of sample size with respect to $d$ and $p$, we assume $\gamma:=\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}=1$ and $\bm{\Sigma}=\bm{I}-{\bm{{\mu}}}{\bm{{\mu}}}^{\sf T}$ the projection onto the orthogonal space of ${\bm{{\mu}}}$. Also note that without loss of generality, we can assume that our estimator $\widehat{\bm{\theta}}$ has unit norm, because its norm does not affect the sign of $\langle\bm{x},{\widehat{\bm{\theta}}}\rangle$. This way, ${\bm{{\mu}}}$ and ${\widehat{\bm{\theta}}}$ both belong to $\mathbb{S}^{d-1}$.
@@
For convenience, we define the misclassification error of a classifier $\widehat{\bm{\theta}}\in\mathbb{S}^{d-1}$ as $\textup{err}(\widehat{\bm{\theta}};{\bm{{\mu}}})=1-\textup{acc}(\widehat{\bm{\theta}};{\bm{{\mu}}})$. As per our condition,
@@
Next recall the set $\mathcal{T}:=\{(\bm{x}_{j},\widehat{y}_{j})\}_{j\in[n]}$, and let ${\bm{X}}=[\bm{x}_{1}^{T},\dotsc,\bm{x}_{n}^{T}]^{T}\in\mathbb{R}^{n\times d}$ and $\widehat{\bm{y}}=[\widehat{y}_{1},\dotsc,\widehat{y}_{n}]^{T}$. By an application of Fano’s inequality, with conditioning on $\mathbf{X}$ (see e.g. (Scarlett and Cevher,, 2019, Section 2.3)) we have
@@
We will now upper bound $\mathcal{I}(I;\widehat{\bm{y}}|{\bm{X}})$. Let $w_{j}:=\frac{1}{2}\big{(}{\rm sign}(\langle\bm{x}_{j},{\bm{{\mu}}}_{I}\rangle)+1\big{)}$ be the 0-1 version of the actual label of $\bm{x}_{j}$, viz., ${\rm sign}(\langle\bm{x}_{j},{\bm{{\mu}}}_{I}\rangle)$. As per our setting, we have $\widehat{y}_{j}=2\big{(}w_{j}\oplus z_{j}\big{)}-1$ where $z_{j}\sim\text{Bernoulli}(p)$ and $\oplus$ denotes modulo-2 addition. Since the noise variables $z_{j}$ are independent and $\widehat{y}_{j}$ depends on $(I,{\bm{X}})$ only through $w_{j}=\frac{1}{2}\big{(}{\rm sign}(\langle\bm{x}_{j},{\bm{{\mu}}}_{I}\rangle)+1\big{)}$, by using the tensorization property of the mutual information (see e.g. (Scarlett and Cevher,, 2019, Lemma 2, part (iii))), we have
@@
Note that $\rho\in(0,1)$ can be set arbitrarily. Since our claim is on the order of sample size, the specific value of $\rho$ does not matter.
We now prove Lemmas D.1, D.2 and D.3.
@@
Note that in eq. D.10 we used the fact that $1+kq_{i}\geq 0$ and $\log x\leq x-1$ for all $x\geq 0$.
@@
We next use Lemma 2.2 from Ball et al., (1997) by which we have $\frac{\mathcal{C}(\bm{v}_{i},\rho)}{\textup{Vol}(\mathbb{S}^{d-1})}\leq e^{-d\rho^{2}/2}$. Using this above, we get
@@
Appendix E Proof of Theorem 4.8
Similar to the proof of Theorem 4.1, we want to upper bound $\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0)$. Plugging in ${\widehat{\bm{\theta}}}_{1}$ from (4.7) we have
@@
A major complication is that ${\widehat{\bm{\theta}}}_{0}$ depends on all of the data points in the training set. Expanding ${\widehat{\bm{\theta}}}_{0}$ we have
${\widehat{\bm{\theta}}}_{0}=\frac{1}{n}\sum_{i}\widehat{y}_{i}\bm{x}_{i}=\frac{1}{n}\sum_{i}\xi_{i}y_{i}\bm{x}_{i}=\frac{1}{n}\left((\sum_{i}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\sum_{i}\xi_{i}y_{i}{\bm{z}}_{i}\right)$ (E.1)
@@
  • For every $\ell\in[n]$ we define the label
    $\widetilde{y}_{\ell}^{\prime}={\rm sign}\left(\langle\bm{x}_{\ell},(\sum_{i}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\right)\,.$ (E.2)
@@
    $\leq\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0;\mathcal{E})+\mathbb{P}(\mathcal{E}^{c})\,.$ (E.3)
@@
    We will control the probability of the RHS above being negative using concentration results on the sum of i.i.d. subgaussian random variables.
Next we get into details of this proof sketch. We start by bounding $\mathbb{P}(\mathcal{E}^{c})$.
Lemma E.1
Fix $\ell\in[n]$. Suppose that $nd\geq\frac{\gamma^{4}}{\lambda_{\max}^{2}}$ and define $p^{\prime}:=(1+\frac{3\gamma^{4}}{8\lambda_{\max}^{2}nd})p$. We then have
$\mathbb{P}\left(\widetilde{y}_{\ell}^{\prime}\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle<0\right)\leq\frac{1}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,.$
Using Lemma E.1 along with union bounding over $\ell\in[n]$ (note that the events are dependent), we get
$\mathbb{P}(\mathcal{E}^{c})\leq\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,.$ (E.4)
@@
$y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle=\frac{y}{n}\langle\bm{x},\sum_{\ell}\widetilde{y}_{\ell}^{\prime}\bm{x}_{\ell}\rangle=\frac{1}{n}\sum_{\ell\in[n]}y{\rm sign}\left(\langle\bm{x}_{\ell},(\sum_{i}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\right)\langle\bm{x},\bm{x}_{\ell}\rangle.$ (E.5)
We define the shorthand $\beta:=\sum_{i}\xi_{i}(1+u_{i})$ and condition on $\beta$. Then (E.5) will be sum of i.i.d. subgaussian variables
$T_{\ell}:=y{\rm sign}\left(\langle\bm{x}_{\ell},\beta{\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\right)\langle\bm{x},\bm{x}_{\ell}\rangle\,,\quad\ell\in[n]\,.$ (E.6)
@@
Lemma E.2
Assuming $\frac{\beta\gamma^{2}}{2\lambda_{\max}d}<1$, for $\ell\in[n]$ we have
@@
where $q=\exp\left(-\frac{\beta\gamma^{2}}{20\lambda_{\max}}\right)$.
@@
$\leq 4|\langle\bm{x},{\bm{{\mu}}}\rangle|+2\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}\,.$
@@
Next by using Hoeffding-type inequality for sum of sub-gaussian random variables (see e.g (Vershynin,, 2010, Proposition 5.1)) we have for every $t\geq 0$,
@@
$\leq 2\exp\left(-\frac{n(1-2q)^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\,,$ (E.7)
where the last step follows from Lemma E.2.
Our next step is to take expectation of the above with respect to $\beta$. Before proceeding we define the following event:
$\mathcal{E}_{0}:=\left\{\beta\geq n(1-2p)/2\right\}\,.$ (E.8)
@@
By choosing $t=n(1-2p)/2$, and using the fact that $u_{i}>0$ for all $i$ we get
@@
We now continue by taking expectation of (E.7) with respect to $\beta$.
@@
$\leq 2\operatorname{\mathbb{E}}\left[\exp\left(-\frac{n(1-2q)^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\mathbb{1}_{\mathcal{E}_{0}}\right]+2\mathbb{P}(\mathcal{E}_{0}^{c})\,.$ (E.9)
@@
Using the above bound along with the bound on $\mathbb{P}(\mathcal{E}_{0}^{c})$ into (E.9), we arrive at
$\mathbb{P}(\sum_{\ell}T_{\ell}<0)\leq 2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$ (E.10)
Invoking (E.5) and the definition of $T_{\ell}$ given by (E.6) we obtain
$\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0;\mathcal{E})\leq 2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$ (E.11)
For the final step we combine (E.11), (E.4) and (E.3) to get
$\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0)\leq$ $2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)$
$+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)+\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,.$ (E.12)
E.1 Proof of Lemma E.1
Define $\tilde{\xi}_{i}:=\xi_{i}(1+u_{i})$. By invoking (E.1) we have
@@
$\stackrel{{\scriptstyle(d)}}{{\leq}}\frac{1}{2}\exp\left(-\frac{\gamma^{4}(\sum_{i}\tilde{\xi}_{i})^{2}}{4n\lambda_{\max}^{2}\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}\right)\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{2n}\right)\,.$ (E.13)
Here, $(a)$ follows by the choice of $\widetilde{y}^{\prime}_{\ell}$; $(b)$ holds by using (Abramowitz and Stegun,, 1968, Equation (7.1.13)):
@@
for all $t\geq 0$. In addition $(c)$ holds since $(a+b)^{2}\geq\frac{a^{2}}{2}-b^{2}$. Also, $(d)$ holds because under our data model we have $\bm{\Sigma}^{1/2}\bm{x}_{\ell}=\bm{\Sigma}\bm{z}_{\ell}$ and so $\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}\leq\lambda_{\max}\left\|\bm{z}_{\ell}\right\|_{\ell_{2}}$. Further, $|\langle\bm{x}_{\ell},{\bm{{\mu}}}\rangle|=|y(1+u_{\ell})\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}|\geq\gamma^{2}$ because $u_{\ell}\geq 0$. Putting these bounds together we have

### v3

@@ B.2 Upper bound $\widetilde{\alpha}_{0}(\mathbf{x})$
${\sf err}({\widehat{\bm{\theta}}}_{0})\leq\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\gamma^{4}}{16\lambda_{\max}^{2}d}\right)+e^{-d/8}+2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$
@@
Proof.
Since we are interested with dependence of sample size with respect to $d$ and $p$, we assume $\gamma:=\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}=1$ and $\bm{\Sigma}=\bm{I}-{\bm{{\mu}}}{\bm{{\mu}}}^{\sf T}$ the projection onto the orthogonal space of ${\bm{{\mu}}}$. Also note that without loss of generality, we can assume that our estimator $\widehat{\bm{\theta}}$ has unit norm, because its norm does not affect the sign of $\langle\bm{x},{\widehat{\bm{\theta}}}\rangle$. This way, ${\bm{{\mu}}}$ and ${\widehat{\bm{\theta}}}$ both belong to $\mathbb{S}^{d-1}$.
@@
For convenience, we define the misclassification error of a classifier $\widehat{\bm{\theta}}\in\mathbb{S}^{d-1}$ when the ground truth separator is ${\bm{{\mu}}}$ as $\textup{err}(\widehat{\bm{\theta}};{\bm{{\mu}}})=1-\textup{acc}(\widehat{\bm{\theta}};{\bm{{\mu}}})$. As per our condition,
@@
Next recall the set $\mathcal{T}:=\{(\bm{x}_{j},\widehat{y}_{j})\}_{j\in[n]}$, and let ${\bm{X}}=[\bm{x}_{1}^{T},\dotsc,\bm{x}_{n}^{T}]^{T}\in\mathbb{R}^{n\times d}$ and $\widehat{\bm{y}}=[\widehat{y}_{1},\dotsc,\widehat{y}_{n}]^{T}$. By an application of Fano’s inequality, with conditioning on $\mathbf{X}$ (see e.g., Section 2.3 of Scarlett & Cevher (2019)) we have
@@
We will now upper bound $\mathcal{I}(I;\widehat{\bm{y}}|{\bm{X}})$. Let $w_{j}:=\frac{1}{2}\big{(}{\rm sign}(\langle\bm{x}_{j},{\bm{{\mu}}}_{I}\rangle)+1\big{)}$ be the 0-1 version of the actual label of $\bm{x}_{j}$, viz., ${\rm sign}(\langle\bm{x}_{j},{\bm{{\mu}}}_{I}\rangle)$. As per our setting, we have $\widehat{y}_{j}=2\big{(}w_{j}\oplus z_{j}\big{)}-1$ where $z_{j}\sim\text{Bernoulli}(p)$ and $\oplus$ denotes modulo-2 addition. Since the noise variables $z_{j}$ are independent and $\widehat{y}_{j}$ depends on $(I,{\bm{X}})$ only through $w_{j}=\frac{1}{2}\big{(}{\rm sign}(\langle\bm{x}_{j},{\bm{{\mu}}}_{I}\rangle)+1\big{)}$, by using the tensorization property of the mutual information (see e.g., Lemma 2, part (iii) of Scarlett & Cevher (2019)), we have
@@
Note that $\rho\in(0,1)$ can be set arbitrarily. Since our claim is on the order of $n$, the specific value of $\rho$ can be chosen based on $\delta$ to ensure $\frac{\rho^{2}}{2}(1-\delta_{0})$ is of the order of $(1-\delta)$. This finishes the proof of Theorem 4.6. ∎
We will now prove Lemmas D.1, D.2 and D.3.
@@
Note that in Equation D.10 we used the fact that $1+kq_{i}\geq 0$ and $\log x\leq x-1$ for all $x\geq 0$.
@@
We next use Lemma 2.2 from Ball et al. (1997) by which we have $\frac{\mathcal{C}(\bm{v}_{i},\rho)}{\textup{Vol}(\mathbb{S}^{d-1})}\leq e^{-d\rho^{2}/2}$. Using this above, we get
@@
Appendix E Comparison with Other Lower Bounds for Learning with Noisy Labels
Cai & Wei (2021); Im & Grigas (2023); Zhu et al. (2024) derive some noteworthy lower bounds in much more general settings than ours; naturally, their bounds are weaker than our bound in Theorem 4.6. In the work of Cai & Wei (2021), our setting corresponds to $n_{p}=0,n_{q}=n$. As per Theorem 3.2 of Cai & Wei (2021), the lower bound on the error is effectively [^9] This is because $\beta\leq 1$ as per Definition 2 and $\alpha\beta\leq d$ as per the paragraph after Remark 3 (of Cai & Wei (2021)). $n^{-O(\frac{1+\alpha}{d})}$. So when $\alpha\ll d$, this lower bound yields a much worse sample complexity than our result in Theorem 4.6. In Im & Grigas (2023), the lower bound on the error (Theorem 2) does not reduce with $n$, so even if there are infinite samples, we cannot get zero error in the worst case. As for Zhu et al. (2024), their lower bound on the error (Theorem 1) also has a non-diminishing term depending on $\epsilon$. In the special case of $\epsilon=0$ (or $\epsilon$ being small enough), there is an $n^{-1/2}$ dependence but no dependence on the dimension or a related quantity. However, the upper bound in Theorem 2 therein does have a dependence on a VC dimension-like quantity as expected, so their lower bound is probably loose w.r.t. dimension.
Appendix F Proof of Theorem 4.8
Similar to the proof of Theorem 4.1, we want to upper bound $\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0)$. [^10] Note that $\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle)\neq y)=\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0)$. Plugging in ${\widehat{\bm{\theta}}}_{1}$ from (4.8) we have
@@ Appendix E Comparison with Other Lower Bounds for Learning with Noisy Labels
A major complication is that ${\widehat{\bm{\theta}}}_{0}$ depends on all the data points in the training set. Expanding ${\widehat{\bm{\theta}}}_{0}$, we get
${\widehat{\bm{\theta}}}_{0}=\frac{1}{n}\sum_{i}\widehat{y}_{i}\bm{x}_{i}=\frac{1}{n}\sum_{i}\xi_{i}y_{i}\bm{x}_{i}=\frac{1}{n}\left((\sum_{i}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\sum_{i}\xi_{i}y_{i}{\bm{z}}_{i}\right).$ (F.1)
Thus,
$\widetilde{y}_{\ell}={\rm sign}(\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle)={\rm sign}\left(\langle\bm{x}_{\ell},(\sum_{i}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\sum_{i}\xi_{i}y_{i}{\bm{z}}_{i}\rangle\right)\,.$ (F.2)
@@
  • For every $\ell\in[n]$ we define the dummy label
    $\widetilde{y}_{\ell}^{\prime}={\rm sign}\left(\langle\bm{x}_{\ell},(\sum_{i}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\right)\,.$ (F.3)
    Note that $\widetilde{y}_{\ell}^{\prime}$ depends only on ${\bm{z}}_{\ell}$, whereas $\widetilde{y}_{\ell}$ depends on all $\{{\bm{z}}_{i}\}_{i\in[n]}$.
@@
    $\leq\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0;\mathcal{E})+\mathbb{P}(\mathcal{E}^{c})\,.$ (F.4)
@@
    We will control the probability of the RHS above being negative using concentration results on the sum of i.i.d. subgaussian random variables (with appropriate conditioning first).
Next we get into details of this proof outline. We start by bounding $\mathbb{P}(\mathcal{E}^{c})$.
Lemma F.1.
Fix $\ell\in[n]$. Suppose that $nd\geq\frac{\gamma^{4}}{\lambda_{\max}^{2}}$. Define $p^{\prime}:=(1+\frac{3\gamma^{4}}{8\lambda_{\max}^{2}nd})p$. We then have
$\mathbb{P}\left(\widetilde{y}^{\prime}_{\ell}\neq{\rm sign}(\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle)\right)=\mathbb{P}\left(\widetilde{y}_{\ell}^{\prime}\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle<0\right)\leq\frac{1}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,.$
Using Lemma F.1 along with union bounding over $\ell\in[n]$ (note that the events are dependent), we get
$\mathbb{P}(\mathcal{E}^{c})\leq\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n},$ (F.5)
where $p^{\prime}=(1+\frac{3\gamma^{4}}{8\lambda_{\max}^{2}nd})p$.
@@
$y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle=\frac{y}{n}\langle\bm{x},\sum_{\ell}\widetilde{y}_{\ell}^{\prime}\bm{x}_{\ell}\rangle=\frac{1}{n}\sum_{\ell\in[n]}y{\rm sign}\left(\langle\bm{x}_{\ell},(\sum_{i}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\right)\langle\bm{x},\bm{x}_{\ell}\rangle.$ (F.6)
We define the shorthand $\beta:=\sum_{i}\xi_{i}(1+u_{i})$ and condition on $\beta$. Then (F.6) will be sum of i.i.d. subgaussian variables
$T_{\ell}:=y{\rm sign}\left(\langle\bm{x}_{\ell},\beta{\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\right)\langle\bm{x},\bm{x}_{\ell}\rangle\,,\quad\ell\in[n]\,.$ (F.7)
@@
Lemma F.2.
Suppose that $\frac{\beta\gamma^{2}}{2\lambda_{\max}d}>1$ and $d\geq 7$. Then for $\ell\in[n]$, we have
@@
where $q:=\exp\left(-\frac{\beta\gamma^{2}}{20\lambda_{\max}}\right)$.
@@
$\leq 4|\langle\bm{x},{\bm{{\mu}}}\rangle|+2\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}},$
@@
Next by using Hoeffding-type inequality for sum of sub-gaussian random variables (see e.g., Proposition 5.1 of Vershynin (2010)) we have for every $t\geq 0$,
@@
$\leq 2\exp\left(-\frac{n(1-2q)^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\,,$ (F.8)
where the last step follows from Lemma F.2. Recall that $q=\exp\left(-\frac{\beta\gamma^{2}}{20\lambda_{\max}}\right)$.
Our next step is to take expectation of the above with respect to $\beta=\sum_{i}\xi_{i}(1+u_{i})$. Before proceeding we define the following event:
$\mathcal{E}_{0}:=\left\{\beta\geq n(1-2p)/2\right\}\,.$ (F.9)
@@
By choosing $t=n(1-2p)/2$, and using the fact that $u_{i}>0$ for all $i$, we get
@@
We now continue by taking expectation of (F.8) with respect to $\beta$.
@@
$\leq 2\operatorname{\mathbb{E}}\left[\exp\left(-\frac{n(1-2q)^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\mathbb{1}_{\mathcal{E}_{0}}\right]+2\mathbb{P}(\mathcal{E}_{0}^{c})\,.$ (F.10)
@@
Using the above bound along with the bound on $\mathbb{P}(\mathcal{E}_{0}^{c})$ into (F.10), we arrive at
$\mathbb{P}(\sum_{\ell}T_{\ell}<0)\leq 2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$ (F.11)
Invoking (F.6) and the definition of $T_{\ell}$ given by (F.7) we obtain
$\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0;\mathcal{E})\leq 2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$ (F.12)
Finally we combine (F.12), (F.5) and (F.4) to get
$\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0)$ $\leq 2\exp\left(-\frac{n(1-2q^{\prime})^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)$
$+4\exp\left(-\frac{n}{32}(1-2p)^{2}\right)+\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}.$ (F.13)
This finishes the proof of Theorem 4.8.
We will now prove the intermediate lemmas F.1 and F.2.
F.1 Proof of Lemma F.1
First note that $\mathbb{P}\left(\widetilde{y}^{\prime}_{\ell}\neq{\rm sign}(\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle)\right)=\mathbb{P}\left(\widetilde{y}_{\ell}^{\prime}\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle<0\right)$.
Define $\tilde{\xi}_{i}:=\xi_{i}(1+u_{i})$. Then using (F.1), we have
@@
$\stackrel{{\scriptstyle(d)}}{{\leq}}\frac{1}{2}\exp\left(-\frac{\gamma^{4}(\sum_{i}\tilde{\xi}_{i})^{2}}{4n\lambda_{\max}^{2}\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}\right)\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{2n}\right)\,.$ (F.14)
Here, $(a)$ follows by the choice of $\widetilde{y}^{\prime}_{\ell}$; $(b)$ holds by using Equation (7.1.13) of Abramowitz & Stegun (1968):
@@
for all $t\geq 0$. In addition $(c)$ holds since $(a+b)^{2}\geq\frac{a^{2}}{2}-b^{2}$. Also, $(d)$ holds because under our data model we have $\bm{\Sigma}^{1/2}\bm{x}_{\ell}=\bm{\Sigma}\bm{z}_{\ell}$ and so $\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}\leq\lambda_{\max}\left\|\bm{z}_{\ell}\right\|_{\ell_{2}}$. Further, $|\langle\bm{x}_{\ell},{\bm{{\mu}}}\rangle|=|y(1+u_{\ell})\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}|\geq\gamma^{2}$ because $u_{\ell}\geq 0$. Putting these bounds together, we have